### v1

# Learning from Reasoning Failures via Synthetic Data Generation

Abstract
Training models on synthetic data has emerged as an increasingly important strategy for improving the performance of generative AI. This approach is particularly helpful for large multimodal models (LMMs) due to the relative scarcity of high-quality paired image-text data compared to language-only data. While a variety of methods have been proposed for generating large multimodal datasets, they do not tailor the synthetic data to address specific deficiencies in the reasoning abilities of LMMs which will be trained with the generated dataset. In contrast, humans often learn in a more efficient manner by seeking out examples related to the types of reasoning where they have failed previously. Inspired by this observation, we propose a new approach for synthetic data generation which is grounded in the analysis of an existing LMM’s reasoning failures. Our methodology leverages frontier models to automatically analyze errors produced by a weaker LMM and propose new examples which can be used to correct the reasoning failure via additional training, which are then further filtered to ensure high quality. We generate a large multimodal instruction tuning dataset containing over 553k examples using our approach and conduct extensive experiments demonstrating its utility for improving the performance of LMMs on multiple downstream tasks. Our results show that models trained on our synthetic data can even exceed the performance of LMMs trained on an equivalent amount of additional real data, demonstrating the high value of generating synthetic data targeted to specific reasoning failure modes in LMMs. We will make our dataset and code publicly available.
1 Introduction
Recent advancements in large language models (LLMs) have significantly advanced the state-of-art across a broad range of problem domains, which can be attributed in large part to scaling the size of models and datasets used for training. However, a limiting factor to further scaling of training datasets is the availability of new high-quality data. With recent studies suggesting that we are rapidly approaching saturation of existing public data sources [39, 42], synthetic data generation has become an increasingly popular approach for augmenting training datasets. The generation of high-quality synthetic data has become more feasible recently as the capabilities of LLMs have increased, thereby enabling their use for the production of synthetic data at scale [7, 45, 20, 13, 14, 35].
Synthetic data has been particularly valuable for training large multimodal models (LMMs), which combine an LLM with a vision encoder to enable text generation conditioned on multimodal inputs. In contrast to LLMs, real data suitable for training LMMs is relatively scarce due to the lack of naturally-occurring images paired with the type of high-quality text expected to be produced by such models. Consequently, popular LMMs such as LLaVA [22] use partially synthetic data for multimodal instruction tuning, which is often collected by generating synthetic text for real images.
Figure 1: Illustration of our approach. Given a sample from an existing dataset which LLaVA answers incorrectly, we prompt a frontier model to analyze LLaVA’s reasoning failures and propose new synthetic samples which require similar types of reasoning.
Existing approaches to generating synthetic data suitable for training LMMs suffer from two primary limitations. First, most rely on the acquisition of real images, which are then paired with synthetic text produced by another LMM. This limits their application to domains where images are readily available and therefore makes such approaches ill-suited for low-resource settings where image data is scarce. Second, prior approaches utilize broad generation strategies which arbitrarily produce synthetic examples without consideration of the type of data which would be most useful for improving LMMs. This results in an inefficient approach for both generating data and training LMMs, as many of the synthetic examples may provide little or no incremental value when added to training datasets.
In contrast to the dominant paradigm for generating synthetic training data, humans often learn more efficiently by focusing their attention on examples which are most related to their past reasoning failures. To acquire expert knowledge, Ericsson et al. [8] argue that humans should “ideally be given explicit instructions about the best method and be supervised by a teacher to allow individualized diagnosis of errors, informative feedback, and remedial part training.” Since humans acquire new knowledge by learning from their failures [6, 3, 49, 30], seeking out additional examples related to past reasoning errors for further practice and evaluation helps the learner achieve expertise in a subject. Problems which require types of reasoning that have already been mastered are often ignored under this paradigm, as they provide little incremental value in the learning process.
Motivated by such observations of human learning behavior, we propose a new approach for generating synthetic multimodal data which is grounded in an existing LMM’s reasoning failures. First, we identify the failure modes of an existing LMM by evaluating it on the training split of a benchmark dataset and then asking a strong frontier model to hypothesize possible reasons for its incorrect predictions. The frontier model then proposes new question & answer pairs related to the identified reasoning failure modes as well as descriptions of images to accompany them, which can be either used to link the new examples to existing images, or as prompts to generate synthetic images. We further filter the resulting synthetic data using an LMM-as-a-judge methodology to ensure high quality.
Using our approach, we generate a large multimodal instruction tuning dataset containing over 553k examples derived from the reasoning failures of LLaVA-1.5-7B, a popular LMM. We conduct extensive training experiments showing that our synthetic data improves LLaVA’s performance on multiple downstream tasks, even outperforming additional training on alternative real datasets. This is notable, as prior studies have shown that training on synthetic data often requires significantly more examples to achieve equivalent results as training on real data [12]. Training on our dataset also produces greater performance gains than previously-proposed synthetic datasets, highlighting the utility of grounding synthetic data generation in the analysis of a model’s reasoning failures.
To summarize, our contributions are as follows:
  1. 1.
    We propose a new approach for generating fully synthetic multimodal datasets which is grounded in an analysis of an existing LMM’s reasoning failures.
  2. 2.
    Using our methodology, we produce a large multimodal instruction tuning dataset containing over 553k examples derived from errors produced by LLaVA.
  3. 3.
    We conduct extensive experiments demonstrating the utility of our dataset for improving LMMs via training.
  4. 4.
    Through ablations and human analyses, we demonstrate the high quality of data produced by our approach.
  5. 5.
    We make our dataset and code publicly available upon publication to support future research.
2 Related Work
Synthetic datasets for training LMMs.
Chen et al. [5] proposed the ALLaVA dataset, which consists of 1.3M samples of real images paired with image annotations and question-answer pairs generated by a frontier LMM. While their dataset is similarly suited for multimodal instruction tuning as our dataset, they utilize only real images and do not ground data generation in model failures, whereas our approach can utilize both real and synthetically generated images. Li et al. [21] generate synthetic question-answer pairs from an LLM for real chart images to address the narrower domain of reasoning-based chart VQA. Yang et al. [44] generate synthetic multimodal data for text-rich images, leveraging code-guided generation in languages such as LaTeX and HTML to automatically generate synthetic images with dense text descriptions. In contrast, our approach can be broadly applied to different problem domains while leveraging text-to-image diffusion models to produce more diverse types of images.
Synthetic data generation from model failures.
Some prior work has explored the use of model failures in the context of synthetic data generation. Jain et al. [15] proposed an approach to identify directions in a vision model’s latent space which correspond to model failures, which can then be integrated into diffusion models to generate synthetic images tailored to correcting the model’s mistakes. Chegini and Feizi [4] utilize ChatGPT and CLIP to identify text descriptions corresponding to a vision model’s failure models, which are used as prompts for diffusion models to produce synthetic images for data augmentation. In the realm of language-only models, DISCERN [28] utilizes two LLMs to iteratively produce natural language descriptions of errors for synthetic data generation. Lee et al. [20] use incorrect answers from a student LLM finetuned on specific tasks as input to a teacher LLM which generates new examples to use for training. These prior works differ from ours in that they (1) generate only single-modality data (e.g., text-only or image-only datasets), and (2) focus on classification tasks such as image recognition or text classification. In contrast, our approach generates multimodal image-text datasets aimed at training models for open-ended text generation conditioned on multimodal inputs.
Generating synthetic data from frontier models to teach new skills.
AgentInstruct [29] is an agentic framework for generating synthetic data from a powerful frontier model (e.g., GPT-4) to teach new skills to a weaker LLM. Similarly, Ziegler et al. [48] utilize few-shot examples annotated by humans and retrieved documents with produce synthetic data from LLMs for teaching specialized tasks to models. Prompt-based methods for synthetic data generation from LLMs without seed documents [10, 37] as well as knowledge distillation from a teacher model [16] have also been proposed. Unlike our work, these prior studies focus on language-only data generation and use seed documents (e.g., raw text, source code) or prompts as a basis for data generation rather than an analysis of model failures.
3 Dataset construction
3.1 Diagnosing model failures
To generate our tailored synthetic data, we first analyze reasoning failures in a baseline LMM using a more advanced frontier LMM. The frontier model is selected for its superior multimodal reasoning capabilities and high accuracy on diverse vision-language benchmarks. Reasoning failures are identified by evaluating both models on the training sets of vision-language benchmarks, and selecting samples where the baseline LMM produces incorrect responses, while the frontier model succeeds. This process generates a subset of failure cases per benchmark, which we denote as Model Failure Sets (MFS). Each MFS highlights the weaknesses of the baseline LMM, providing a focused challenging dataset.
3.2 Synthetic data generation
Using the resulting MFS, we guide the frontier model through a structured multi-turn process to diagnose the failure reasons of the baseline LMM and generate new training samples that address these failure modes, as illustrated in Figure 5. In the first steps, the frontier model is prompted to describe the image and analyze reasoning errors by examining the question, ground truth answer, and the incorrect response generated by the baseline LMM. Next, the frontier model is instructed to propose new challenging samples which consist of a detailed image description, a clear question, and a deterministic answer designed to target the identified failure modes. We explore two different approaches for sourcing images corresponding to these samples: one which utilizes existing real images, and another which uses synthetically generated images.
Method 1: Question-answer generation for existing images.
We leverage the original images from failed samples and prompt the frontier model to generate 10 new question-answer pairs per sample using the reasoning pipeline described above. This approach is particularly effective for datasets that require precise text rendering or spatial accuracy, such as InfoVQA and ScienceQA. For these types of benchmarks, text-to-image models often struggle to generate coherent images that accurately reflect the textual content or capture the reasoning embedded in the prompt. This method utilizes the prompt described in Figure 5 (Section 7.3 of the supplementary material), but with the final step omitted.
Method 2: Fully synthetic question, answer, and image generation.
In this approach, the frontier model generates both a question-answer pair and a detailed image prompt, which is used to prompt a text-to-image diffusion model to generate synthetic images. For each failed sample, we create 100 fully synthetic samples, consisting of 10 (image prompt, question, answer) triplets, with each image prompt generating 10 images using different classifier-free guidance scales to enhance image diversity. This method utilizes the complete prompt depicted in Figure 5 (Section 7.3 of the supplementary material).
Encouraging diversity.
To further enhance the diversity of our data, we use a variation of our prompt which instructs the frontier LMM to “provide examples that challenge Model A’s weaknesses using scenarios from entirely different domains or situations than the original example.” This additional instruction is appended to Step 4 (Figure 5) to encourage the generation of samples in a different domain. Figure 2 compares fully synthetic domain similar and non-similar samples. It is notable that domain-similar samples preserve the original theme, while non-similar samples cover a more diverse contextual range to improve generalization. Additionally, we created samples where we both enforced and relaxed constraints on question format (e.g., multiple-choice, true/false) and instructions (e.g., requiring responses like ”Unanswerable” when information was insufficient or limiting answers to short responses, see the Shiba Inu example from Figure 9).
Figure 2: Comparison of fully synthetic similar and non-similar samples. Similar samples maintain a children’s characters-based theme like the original sample, while non-similar samples address the failure modes by introducing diverse contexts.
3.3 Filtering
We apply a filtering process using the same frontier LMM which produced the samples. The LMM is instructed to evaluate each synthetic sample given an image, question, and answer on a scale of 1 to 3, where 1 indicates an incorrect sample, 2 being partially correct, and 3 fully correct. The prompt used for filtering is provided in Figure 6 of Section  7.4. Only samples rated 3 were included in the final dataset to ensure quality and reliability.
Figure 3: Examples of generated synthetic question-answer pairs for real images from VizWiz, InfoVQA, and ScienceQA.
3.4 Details of generated dataset
Our synthetic dataset consists of 553,992 samples incorporating both real and generated images derived from the MFS of LLaVA-1.5-7B on four benchmark training sets: VizWiz [11], InfoVQA [27], ScienceQA [25], and OK-VQA [26]. These benchmarks were chosen to ensure a wide range of visual and reasoning challenges. VizWiz consists of real-world images captured by visually impaired users, often requiring detailed scene understanding. OK-VQA focuses on visual question answering which requires external knowledge. InfoVQA involves text-rich images where reading comprehension is crucial, assessing the model’s ability to extract and interpret textual information from images. ScienceQA includes multimodal scientific reasoning questions which require both spatial and logical reasoning, making it valuable for evaluating complex reasoning capabilities. We use GPT-4o [31] as the frontier LMM for analyzing the reasoning failures of LLaVA-1.5-7B due to its strong multimodal reasoning capabilities. To generate the synthetic images, we utilized FLUX.1-schnell Labs [18] text-to-image model with a resolution of 1024×1024 pixels and a classifier guidance scale ranging from 3 to 13.
Dataset LLAVA $\textrm{LLaVA}_{\textrm{syn}}$
CIFAR-10  [17] 82.1 81.2
Food-101  [2] 13.4 13.2
iNaturalist  [38] 20.6 52.0
MNIST  [19] 75.1 80.5
F-MNIST  [41] 9.8 10.0
Oxford-pets  [32] 39.6 96.4
Table 1: Image classification accuracy of LLaVA and a $\textrm{LLaVA}_{\textrm{syn}}$ model augmented only with synthetic examples corresponding to object recognition failures.
Table 5 (Section 7.2 of the supplementary material) provides an overview of each benchmark, including the original size of the training split, the MFS size of LLaVA-1.5-7b, and the total number of synthetic samples retained after filtering. As described in Section 3.2, we utilized two methods for synthetic data generation, depending on the dataset: for InfoVQA and ScienceQA, we applied Method 1, which generates new question-answer pairs based on the original real images. For OK-VQA, we employed Method 2, generating synthetic images along with corresponding question-answer pairs. For VizWiz dataset we integrated both methods, generating additional synthetic images as well as new question-answer pairs based on its existing real images. Overall, our final dataset consists of 42% real-image-based samples and 58% fully synthetic samples. All samples in the final dataset contain a single-turn conversation, featuring a question and a short answer.
Our filtering approach effectively removes lower-quality synthetic samples. For VizWiz, approximately 81% of synthetic-image samples and 34% of real-image-based samples were removed. This indicates, as expected, that generating entirely synthetic samples is more challenging than generating synthetic text alone for real images. OK-VQA had a lower removal rate of 29% for synthetic images compared to VizWiz, suggesting better synthetic image generation quality, possibly resulting from simpler and less ambiguous visual content. Among real-image-based samples, ScienceQA experienced a similar removal rate (29%), likely due to the complexity of spatial and scientific reasoning tasks. In contrast, InfoVQA exhibited a significantly lower removal rate of only 5% with an average filtering score of 2.9 (out of 3), indicating the strong capability of GPT-4o in handling text-based images. Nevertheless, the large number of generated data that passed filtering ensures comprehensive coverage of the reasoning failures previously identified in LLaVA-1.5-7b across various benchmarks. Figures 3 show sampled examples from VizWiz and InfoVQA highlighting the diversity of question types and demonstrating the overall quality of generated images and text. Appendix 10 provides additional examples from our dataset, including instances that are incorrect or ambiguous.
4 Experiments
4.1 Details of experimental setting
We perform vision-language instruction tuning with LLaVA-v1.5-7B, building upon the LLaVA-1.5-mix-665K instruction tuning dataset [23]. This dataset contains 665K structured user-GPT conversations mostly focused on visual prompts. Because our approach explicitly targets image-grounded reasoning, we only selected the 624K samples that include images from the original LLaVA-instruct data. We then augmented this visually oriented subset with our generated synthetic MFS (Section 3.4), varying the number of synthetic samples used according to the requirements of each experiment. The resulting training set combines real image-text conversations and generated synthetic samples that address specific reasoning failures.
For fine-tuning, we used Vicuna-v1.5-7B [47] weights as the LLM backbone and leveraged the pretrained multimodal projector from LLaVA-1.5-7b. We followed the original LLaVA training procedure, ensuring a fair comparison to existing methods (see hyperparameters definition in Appendix 7.5). We refer to models trained with a mixture of the original LLaVA-Instruct dataset and our synthetic data as $\textrm{LLaVA}_{\textrm{syn}}$. As a baseline, we report the performance of LLaVA trained under the same setting but without any additional synthetic data added to the training dataset (i.e., $N_{syn}=0$). Additionally, for each dataset from which reasoning failures were derived for synthetic data generation, we report the performance of a LLaVA model trained on an equivalent amount of real data sourced from the corresponding training dataset (denoted as $\textrm{LLaVA}_{\textrm{real}}$). This provides a measure of the efficiency of our synthetic data relative to training on real in-domain data.
We used a variety of multimodal reasoning benchmarks for evaluating models. For in-domain evaluations (i.e., evaluating on a withheld validation set corresponding to the training set from which reasoning failures were derived), we used the InfoVQA, OK-VQA and ScienceQA validation sets. Since InfoVQA and ScienceQA questions rely heavily on reading text contained in the images, we further evaluated models on TextVQA [34] and OCR-Bench [9]. Finally, because our synthetic dataset was designed to enhance the model’s reasoning capabilities, we chose MMBench [24] and MMMU [46] as additional OOD benchmarks.
4.2 Synthetic data augmentation results
Table 2 provides in-domain evaluation results utilizing synthetic data derived from InfoVQA, ScienceQA, and OK-VQA. Notably, augmenting the LLaVA-Instruct dataset with our synthetic data achieves performance comparable to or better than using an equivalent amount of real domain-specific data in most cases. This result is particularly significant given that the synthetic samples were generated using only a small subset of the original training data: specifically, only those examples where LLaVA scored 0.0 while GPT scored 1.0. For instance, with OK-VQA, the original training set consists of 9,009 samples, but we utilized only 607 training samples which LLaVA failed on in order to generate 9009 synthetic samples, resulting in a performance boost of 13% on the OK-VQA test set. Similarly, our approach utilized only 28% of the ScienceQA training dataset to generate full synthetic replacements, yet still resulted in better performance than training directly on the real dataset. We also observe that performance improves as the amount of synthetic data used for data augmentation increases.
In practice, it may be desirable to combine synthetically generated data which was derived from reasoning failures across different datasets. We therefore provide results for two different sized mixtures of our synthetic data derived from InfoVQA, ScienceQA, and OK-VQA reasoning failures in Table 3. In addition to in-domain evaluations for these three datasets, we provide results for the five other datasets mentioned previously to measure the impact on OOD generalization. We also provide results for LLaVA models trained on three alternative synthetically generated datasets: ALLaVA, CoSyn-400k, and SimVQA [^1] ALLaVA consists of real images with synthetic text, whereas Cosyn-400k and SimVQA contain fully synthetic image-text examples. Our synthetic data utilized here contains both real and synthetic images.. From Table 3, we observe that our synthetic data outperforms almost all baselines (except VQAv2) in the maximum data augmentation setting ($N_{syn}=124,922$). When augmenting training sets with half as much data, we observe that our synthetic data produces the greatest improvements in-domain, whereas the ALLaVA dataset, which contains only real images paired with synthetic text, performs slightly better in other OOD settings. Finally, we note that training LLaVA in this mixed synthetic dataset setting produces similar in-domain performance as training LLaVA separately for reasoning failures derived from each individual dataset (Table 2).
Dataset Model $N$ $N_{syn}$ EM Score
InfoVQA LLaVA 624,610 0 26.7
$\textrm{LLaVA}_{\textrm{real}}$ 634,684 0 31.6
$\textrm{LLaVA}_{\textrm{syn}}$ 634,684 10,074 30.8
$\textrm{LLaVA}_{\textrm{syn}}$ 687,071 62,461 33.0
$\textrm{LLaVA}_{\textrm{syn}}$ 710,610 86,000 34.3
ScienceQA LLaVA 624,610 0 70.7
$\textrm{LLaVA}_{\textrm{real}}$ 630,195 0 70.0
$\textrm{LLaVA}_{\textrm{syn}}$ 630,195 5,585 71.8
$\textrm{LLaVA}_{\textrm{syn}}$ 646,594 21,984 73.0
OK-VQA LLaVA 624,610 0 57.0
$\textrm{LLaVA}_{\textrm{real}}$ 633,619 0 54.3
$\textrm{LLaVA}_{\textrm{syn}}$ 633,619 9,009 61.3
$\textrm{LLaVA}_{\textrm{syn}}$ 687,071 62,461 61.3
$\textrm{LLaVA}_{\textrm{syn}}$ 749,532 124,922 61.5
Table 2: In-domain evaluation of baseline LLaVA, LLaVA models trained using synthetically augmented data ($\textrm{LLaVA}_{\textrm{syn}}$) and using training data augmented with real in-domain data ($\textrm{LLaVA}_{\textrm{real}}$).
Base LLM $N$ $N_{syn}$ Augmentation Data TextVQA OCR-Bench InfoVQA OK-VQA VQAv2 ScienceQA MMBench MMMU
Vicuna-7B 624,610 0 N/A (baseline) 47.0 31.9 26.7 57.0 71.5 70.7 52.3 36.4
Vicuna-7B 687,071 62,461 ALLaVA 47.9 34.0 28.4 50.4 72.6 71.2 50.2 36.7
CoSyn-400K 47.1 31.8 28.5 55.8 72.3 71.5 52.4 34.6
SimVQA 46.8 31.6 26.6 54.4 71.8 71.1 53.5 34.7
Ours 47.4 33.2 33.1 60.8 71.3 73.1 52.5 36.2
Vicuna-7B 749,532 124,922 ALLaVA 47.2 34.1 28.8 49.4 72.8 66.5 43.5 34.2
CoSyn-400K 46.8 32.7 29.6 57.7 72.0 70.9 51.9 36.9
Ours 47.4 34.5 33.2 61.1 72.4 73.0 52.5 37.4
Gemma-2B 624,610 0 Ours 39.9 28.3 21.8 51.7 68.7 62.3 29.2 32.3
749,532 124,922 40.9 29.9 29.8 54.8 69.0 65.2 30.7 31.7
Qwen2-7B 624,610 0 Ours 45 31.4 26.7 59.2 71.7 77.9 63 42.3
749,532 124,922 46.2 32.5 27 60.6 72.9 80.7 63.5 42.2
Table 3: Training data augmentation experimental results. $N$ denotes the total number of training examples, $N_{syn}$ denotes the number of synthetic examples in the training dataset generated using our approach. The first section of the table compared the same base LLM (Vicuna) trained on various datasets with our dataset, while the second section compares different LLM backbones trained on our dataset.
To evaluate the generalization ability of our synthetic dataset, we trained models using the same datasets but with different backbone LLMs. In the following experiments, we used Gemma-2B [36] and Qwen2-7B [43] as base LLMs. We adopted the LLaVA two-phase training procedure: pretraining on the LLaVA 558k dataset [23] followed by instruction fine-tuning. Table 3 indicates that our synthetic data in its maximum augmentation setting, $N_{syn}=124,922$, outperform the baseline for both LLaVA-Gemma-2B and LLaVA-Qwen-7B almost on all benchmarks (except MMMU) Notably, despite being generated based on LLaVA-Vicuna-7B failure mechanisms, our synthetic data enhance the performance of other models whether they are of same size (Qwen2-7B) or smaller (Gemma-2B).
4.3 Correcting specific types of reasoning failures
Figure 4: Figure shows the clusters of LLAVA reasoning failures described by GPT-4o.
Our synthetic data generation approach explicitly identifies different types of LMM reasoning failures. To systematically categorize these failures, we encoded each reasoning explanation using sentence transformers  [33] and clustered them using k-means. Figure  4 presents the resulting clusters, highlighting prevalent failure modes such as optical character recognition (OCR) and object detection errors. Based on this analysis, we further investigated whether targeted synthetic data can effectively address these specific failure cases and enhance LLAVA’s reasoning capabilities.
Specifically, we augmented LLaVA-Instruct with 10,579 synthetic samples from our $\textrm{VizWiz}_{\textrm{syn}}$-MFS addressing object detection reasoning failures and repeated the second stage of LLaVA finetuning. The model was then evaluated on CIFAR-10  [17], Food-101  [2], iNaturalist  [38], MNIST  [19], Fashion-MNIST  [41], and Oxford-Pets (Binary)  [32] by formatting samples as multiple-choice questions. Table  1 presents a comparison of LLaVA and $\textrm{LLaVA}_{\textrm{syn}}$. The results show that $\textrm{LLaVA}_{\textrm{syn}}$ surpasses LLaVA on four out of six datasets, with particularly notable improvements on iNaturalist, MNIST and Oxford-Pets. This demonstrates the significant impact of our synthetic dataset in addressing specific reasoning failures within LLAVA. By systematically incorporating targeted synthetic samples, we can mitigate common failure cases, leading to measurable performance improvements across multiple benchmarks. Our findings highlight the effectiveness of leveraging targeted synthetic data to refine model reasoning and suggest that incorporating such data-driven interventions can significantly enhance the robustness and generalization of LMMs.
5 Analysis
5.1 Human evaluation of dataset quality
Three of the authors of this work conducted a human evaluation by assessing three different aspects of our generated samples: (1) the alignment of the question and answer in relation to the image prompt, (2) the alignment between the image prompt and the generated image, and (3) the correctness of the answer given the question and image. The first evaluation reflects the quality of reasoning, the second evaluates the fidelity of the image generator’s output, and the third combines both aspects. Scores range from 1 to 3, where 1 indicates an irrelevant alignment, 3 signifies a relevant alignment, and 2 represents a partially relevant or ambiguous alignment. We evaluated 200 samples in total, with 101 containing real images and 99 being fully synthetic. The overall correctness score for answers was 2.78, with real‐image‐based samples scoring 2.75 and fully synthetic samples scoring 2.81, indicating that fully synthetic samples achieve a level of fidelity equal to or even slightly exceeding that of real‐image-based samples. For the synthetic samples specifically, we also measured the alignment between the image prompt and the generated image (2.66), and the alignment of the generated question and answer with the image prompt (2.84), indicating the high quality of reasoning in the generated responses.
5.2 Training data substitution vs. augmentation and impact of synthetically generated images
Train Data $N$ $N_{syn}$ TextVQA OCR-Bench InfoVQA VQAv2 OK-VQA ScienceQA MMBench MMMU
Baseline 624,610 0 47.0 31.9 26.7 71.5 57.0 70.7 52.3 36.4
Substitute w/ syn images 624,610 62,461 47.1 31.6 26.5 72.9 56.9 70.8 52.3 35.3
624,610 156153 46.9 31.1 27.0 72.7 57.0 70.6 51.2 37.9
Augment w/ syn images 645,222 20,612 46.7 32.0 27.0 72.6 57.4 71.2 53.4 34.9
687,071 62,461 47.7 31.8 25.8 71.6 59.4 71.2 52.3 36.4
Augment w/ real images 645,222 20,612 46.9 32.5 27.2 72.6 57.4 70.6 53.1 35.0
687,071 62,461 47.2 32.2 27.2 71.8 56.9 71.2 52.3 33.8
Table 4: Ablation experiments comparing baseline LLaVA to LLaVA models trained with synthetic data generated from VizSiz failures. We investigate substitution and augmentation strategies for synthetic data, as well as the use of synthetic vs. real images.
Our previous experiments augmented an existing 624k sample training dataset (LLaVA-Instruct) with our synthetic data. In domains where data is scarce, training datasets of this size may not be available. To investigate the utility of our synthetic data in such low-resource settings, we conducted experiments in which we randomly substituted different quantities of examples from the original dataset with our synthetically generated data [^2] We used synthetic data derived from Vizwiz failures in this setting.. The results of this experiment are provided in rows 2-3 of Table 4. Even when up to 25% of the original dataset is substituted with our synthetic data, we achieve performance that is either as good or better than the baseline LLaVA model across a broad range of downstream tasks. This is despite the fact that the original LLaVA training dataset utilizes real images, whereas our synthetic data used in this experiment contained only synthetically generated images. The fact that our synthetic data achieves similar or better performance than an existing real data source is significant, as prior studies have shown that training on synthetically generated image data is often much less efficient than training on an equivalent amount of real image data [12]. Table 4 also shows the impact of using real vs. synthetic images in our pipeline. Specifically, we compare the effectiveness of our synthetic data derived from Vizwiz reasoning failures when paired with real images (from Vizwiz) or synthetically generated images. In the training data augmentation setting, we observe that synthetic images generally achieve similar results as utilizing real images. Synthetic images even surpass the performance of real images in TextVQA, OK-VQA, and MMBench. This demonstrates the high quality of our synthetic images and their potential to serve as replacements for real images in low-resource settings where data is scarce.
5.3 Impact of filtering on data quality
To investigate the impact of filtering on the quality of synthetically generated data, we repeated our in-domain evaluation experiments for ScienceQA and OK-VQA using raw unfiltered data. In the maximum synthetic data augmentation setting (last row of each section in Table 2), using unfiltered data reduces EM from 73.0 to 72.2 on ScienceQA and from 63.3 to 58.8 on OK-VQA. This shows that our filtering approach improves model performance when using our synthetic examples for training data augmentation. Furthermore, using only synthetic examples which were assigned the lowest rating in our filtering process decreases the EM score on OK-VQA to 57.5, which highlights the difference in quality between the lowest-scoring and highest-scoring synthetic examples identified during filtering.
5.4 Comparison of LLM synthetic data generators
We compared two frontier LMMs, GPT-4o and Qwen2-VL-7B [40], for generating synthetic data grounded in LLaVA-7B failures. Qwen2-VL-7B was chosen for its high accuracy on vision-language benchmarks. However, samples generated by Qwen2-VL reduced downstream performance by 2% on InfoVQA and 6.5% on OK-VQA compared to GPT-4o. Additionally, those samples received lower filtering scores: 1.9 (Qwen2-VL) vs. 2.5 (GPT-4o) for OK-VQA, and 2.6 vs. 2.9 for InfoVQA. Based on our manual analysis, we hypothesiize that these differences may result from the detailed and precise reasoning provided by GPT-4o, resulting in synthetic samples that are better tailored to address identified reasoning failures. In contrast, samples generated by Qwen2-VL-7B sometimes demonstrate lower diversity, which could limit their effectiveness in addressing the broad range of failure modes. Figure 7 (Section 9.1 of the supplementary material) provides an example of these observed differences in the reasoning processes of GPT-4o and Qwen2-VL-7B models, as well as the corresponding generated fully synthetic samples.
6 Conclusion
We developed a novel method for generating multimodal synthetic data by analyzing model reasoning failures. This approach enabled us to create a multimodal instruction tuning dataset with over 553k synthetic examples derived from LLaVA’s errors. Experimental results show that our synthetic data significantly improves LLaVA’s in-domain performance on InfoVQA, ScienceQA, and OK-VQA, even outperforming training on an equivalent amount of real data sampled from most of these datasets. Furthermore, models trained on our synthetic dataset exhibit improvements in OOD evaluations and outperform training on other existing synthetic datasets when the amount of training data augmentation is scaled. We also showed that training LLaVA only on examples derived from specific failure modes improves its performance on tasks which require corresponding forms of reasoning. Through additional ablations and human evaluations, we demonstrated the effectiveness of different components of our methodology and the high quality of our dataset. We believe our study points to the promise of targeted synthetic data generation strategies which leverage an understanding of a model’s reasoning deficiencies to construct better synthetic examples.
Supplementary Material
7 Implementation details
7.1 Compute Infrastructure
To generate our dataset, we queried GPT-4o through the Azure OpenAI API and deployed Qwen2-VL on Nvidia RTX A6000 GPUs. Using Intel® Gaudi 2 AI accelerators from the Intel® Tiber AI Cloud, we generated 1.024 million images from the VizWiz failed samples and 535k images derived from OK-VQA.
7.2 Dataset statistics
Table 5 provides statistics detailing the quantity of synthetic examples in our dataset which were derived from reasoning failures on different benchmarks. Additional discussion of the dataset composition is provided in Section 3.4.
$\textrm{Dataset}_{\textrm{Image Type}}$ Original Failures Filtered
$\textrm{VizWiz}_{\textrm{real}}$ 20,523 7,785 100,280
$\textrm{VizWiz}_{\textrm{syn}}$ 20,523 7,785 190,172
$\textrm{InfoVQA}_{\textrm{real}}$ 10,074 5,250 95,783
$\textrm{ScienceQA}_{\textrm{real}}$ 5,585 1,562 39,090
[…]
Table 5: Dataset statistics across benchmarks, including original training set size, number of failure samples (LLaVA-1.5-7b: 0, GPT-4o: 1), and synthetic samples with filtering score 3.
7.3 Data generation prompt
Figure 5 shows the prompt used to generate fully synthetic question-answer, and image samples based on the failure modes of an LMM, according to Method 2.
You are analyzing the performance of a vision-language model (called Model A). Model A’s answer could deviate from the ground truth due to limitations in visual understanding, interpretation, or reasoning. Step 1: Describe the image. Step 2: Given a question, the Ground truth answer, and Model A’s generated answer, describe any key visual elements that might influence Model A’s interpretation. Step 3: Analyze the reasoning steps Model A might have used to generate its answer, considering both the visual and textual information. Identify any weaknesses, errors, or gaps in Model A response compared to the ground truth. Step 4: Suggest 10 additional challenging detailed examples to address these limitations. Step 5: Transform each example into a detailed prompt designed to generate a clear and realistic image using a text-to-image generation model.
Figure 5: Prompt used to generate fully synthetic image-text samples based on the failure modes of an LMM (Method 2).
7.4 Filtering prompt
Figure 6 provides the prompt which we used for the filtering stage of our synthetic data generation pipeline. See Section 3.3 of the main paper for additional filtering details.
Given sample containing an image, a question, and an answer, your task is to grade the sample from 1 to 3 based on the following criteria: Score 1: The answer is incorrect. Score 2: The answer is correct, but it is one of several possible valid answers. Score 3: The answer is correct, specific, and the only valid answer. The image provides all the necessary context for the answer.
Figure 6: Filtering prompt
7.5 Training hyperparameters
To train our model, we used 8 Nvidia RTX A6000 GPUs using the hyperparameters from Table 6. We employed DeepSpeed ZeRO stage 3 [1] for distributed training.
Batch Size/GPU $16$
Number of GPUs $8$
Gradient Accumulation $1$
Number of epochs $1$
LLaVA Image Size $576$
Optimizer AdamW
Learning Rate $52e-$ 5
BF16 $True$
LR scheduler $cosine$
Vision Tower openai/clip-vit-large-patch14-336
Language Model lmsys/vicuna-7b-v1.5
Table 6: Hyperparameters to train our model.
8 Additional experimental results
8.1 Detailed OOD results for models fit to different subsets of synthetically generated data
Train Dataset $N$ $N_{syn}$ TextVQA OCR-Bench InfoVQA VQAv2 OK-VQA ScienceQA MMBench MMMU
Baseline 624,610 0 0.47 0.32 0.27 0.71 0.57 0.71 52.30 0.36
Vizwiz 645,133 0 0.47 0.28 0.26 0.72 0.59 0.71 51.74 0.38
687,071 62,461 0.48 0.32 0.26 0.72 0.59 0.71 52.25 0.36
749,532 124,922 0.47 0.32 0.27 0.71 0.59 0.70 53.02 0.37
InfoVQA 634,684 0 0.47 0.32 0.32 0.73 0.58 0.70 52.50 0.37
634,684 10,074 0.47 0.33 0.31 0.73 0.59 0.70 52.16 0.36
687,071 62,461 0.47 0.34 0.33 0.71 0.57 0.71 52.69 0.37
710,610 86,000 0.48 0.33 0.34 0.72 0.56 0.71 52.53 0.38
ScienceQA 630,195 0 0.47 0.29 0.26 0.73 0.58 0.70 52.88 0.36
630,195 5,585 0.47 0.32 0.27 0.72 0.57 0.72 53.19 0.37
646,594 21,984 0.47 0.32 0.26 0.72 0.56 0.73 53.12 0.38
OK-VQA 633,619 0 0.47 0.30 0.27 0.72 0.54 0.71 53.35 0.36
633,619 9,009 0.47 0.33 0.28 0.71 0.61 0.71 52.68 0.35
687,071 62,461 0.47 0.33 0.27 0.72 0.61 0.71 51.96 0.35
Table 7: Training data augmentation experimental results. $N$ denotes the total number of examples used for training, while $N_{syn}$ denotes the number of synthetic examples in the training dataset which were generated using our approach.
Table 7 provides additional evaluation results for models trained individually on real and synthetic data derived from Vizwiz, InfoVQA, ScienceQA, and OK-VQA. All reported values are the official evaluation metrics corresponding to each dataset. The first two rows of each section in Table 7 provide a direct comparison of the efficiency of our synthetic data to real data; we observe that augmenting the LLaVA-Instruct dataset with our synthetic data achieves as good or better performance across most settings as augmenting with real domain-specific data. Furthermore, significant performance gains are achieved relative to the LLaVA baseline when our synthetic data is derived from a dataset in the same domain as the benchmark. For example, synthetic data generated from reasoning failures on InfoVQA significantly improve LLaVA’s performance on tasks which require fine-grained text understanding such as OCR-Bench and InfoVQA.
8.2 Generalization to other models
To evaluate the generalization ability of our synthetic dataset, we trained models using the same datasets but with different backbone LLMs. In the following experiments, we used Gemma-2B [36] and Qwen2-7B [43] as base LLMs. We adopted the LLaVA two-phase training procedure: pretraining on the LLaVA 558k dataset followed by instruction fine-tuning. Table 3 indicates that our synthetic data in its maximum augmentation setting, $N_{syn}=124,922$, outperform the baseline for both LLaVA-Gemma-2B and LLaVA-Qwen-7B almost on all benchmarks.
Model $N$ $N_{syn}$ TextVQA OCR-Bench InfoVQA VQAv2 OK-VQA ScienceQA MMBench MMMU
LLaVA-Gemma-2B 624,610 0 0.3992 0.283 0.2181 0.6866 0.5172 0.6227 29.1752 0.3233
62,461 0.4039 0.297 0.292 0.6992 0.5536 0.6484 28.8542 0.3144
124,922 0.4092 0.299 0.2981 0.6902 0.5477 0.6517 30.7322 0.3167
LLaVA-Qwen2-7B 624,610 0 0.4504 0.314 0.2674 0.7168 0.592 0.7793 62.9986 0.4233
62,461 0.4586 0.324 0.2703 0.7208 0.6136 0.7956 62.9900 0.4189
124,922 0.4617 0.325 0.2699 0.7292 0.6064 0.8074 63.4712 0.4222
Table 8: Training data augmentation experimental results. $N$ denotes the total number of examples used for training, while $N_{syn}$ denotes the number of synthetic examples in the training dataset which were generated using our approach. Best result is highlighted in bold and second best result is underlined
9 Additional analyses
9.1 Detailed example comparing different LLMs used for synthetic data generation
Figure 7 provides a detailed example comparing the reasoning and resulting synthetic examples produced by GPT-4o and Qwen2-VL. See Section 5.4 for additional discussion.
Figure 7: Comparison of GPT-4o and Qwen2-VL for generating Failure-Grounded Synthetic Datasets: GPT-4o demonstrates stronger reasoning capabilities, identifying multiple reasoning failures such as missed contextual cues and a lack of correlation between visual elements and the question. In contrast, while Qwen2-VL correctly answering the original question, identifies fewer failure modes and is less accurate in diagnosing LLaVA’s reasoning failures, sometimes focusing on less relevant aspects, such as the kite in the sky. As a result, Qwen2-VL’s generated samples are less diverse, often repeating the same question, whereas GPT-4o’s samples provide broader coverage of identified reasoning failures. Note: GPT-4o’s reasoning is 2–3 times longer than Qwen2-VL’s; only a portion of GPT-4o’s reasoning is shown here, while Qwen2-VL’s reasoning is presented in full.
10 Examples from our dataset
In this section, we present examples from our dataset and highlight its weaknesses and limitations. Figure 9 shows our synthetic data generated from the OK-VQA dataset while Figure 10 corresponds to the VizWiz dataset. In these examples, both the image, the question and the answer are all synthetic. Figure 8 provides additional fully synthetic text & image examples derived from VizWiz and OK-VQA. Figure 11, 12 and 13 illustrate examples derived from the ScienceQA, InfoVQA and VizWiz benchmarks respectively, where the images are real but the questions and answers are synthetically generated. Lastly, Figures 14 and 15 show some incorrect examples for each benchmark.
Figure 8: Examples of fully synthetic samples, using Method 2 as described in 3.2, both question-answer pairs and images were generated.
Figure 9: Examples of generated samples from OK-VQA with synthetic images and synthetic text.
Figure 10: Examples of generated samples from VizWiz with synthetic images and synthetic text.
Figure 11: Examples of generated samples from ScienceQA with real images and synthetic text.
Figure 12: Examples of generated samples from InfoVQA with real images and synthetic text.
Figure 13: Examples of generated samples from VizWiz with real images and synthetic text.
Figure 14: Examples from our dataset real image-synthetic text, where the sample is ambiguous or incorrect.
Figure 15: Examples from our dataset synthetic image-synthetic text, where the sample is ambiguous or incorrect. For readability, the ScienceQA image was cropped to focus on the region of interest.